OpenStudy (anonymous):

how do i evaluate sec(pi/3)? cot (pi/6) and csc(pi/3)? please explain it to me i have a quiz on it soon!

OpenStudy (anonymous):

\[\sec(\theta) = \frac{1}{\cos(\theta)}\] So \[\sec(\frac{\pi}{3}) = \frac{1}{\cos(\frac{\pi}{3})}\]

OpenStudy (anonymous):

Which is \[\frac{1}{\frac{1}{2}}\]. So \[\sec(\frac{\pi}{3}) = 2\]

OpenStudy (anonymous):

\[\cot(\theta) = \frac{1}{\tan(\theta)}\] \[\cot(\frac{\pi}{6}) = \frac{1}{\tan(\frac{\pi}{6})}\] That equals \[\frac{1}{\frac{\sqrt{3}}{3}}\] Which is \[\frac{3\sqrt{3}}{3}\] or just \[\sqrt{3}\]

OpenStudy (anonymous):

Kinda getting the idea??